OpenStudy (anonymous):

25. Tasty Bakery sells three kinds of muffins: chocolate chip muffins for $0.50 each, oatmeal muffins for $0.65 each, and blueberry muffins for $0.75 cents each. Charles buys some of each kind and chooses three more chocolate chip muffins than blueberry muffins. If he spends $8.45 on 14 muffins, how many of each type of muffin does he buy? Write and solve a system of three equations in three variables. Show your work. (4 points)

OpenStudy (aripotta):

0.5x + 0.65y + 0.75z = 8.45 x + y + z = 14 x = z + 3

OpenStudy (aripotta):

do you know what to do?

OpenStudy (aripotta):

@torimoore18 :l

OpenStudy (anonymous):

i understnad the equation but I don't know how to solve it it. Does that make sense?/

OpenStudy (aripotta):

substitute for x in the first two equations

OpenStudy (anonymous):

so guess what x is?

OpenStudy (aripotta):

no, x = z + 3

OpenStudy (anonymous):

i'm still confused..

OpenStudy (aripotta):

i'll do the first equation for ye

OpenStudy (anonymous):

thanks you

OpenStudy (aripotta):

0.5x + 0.65y + 0.75z = 8.45 0.5(z + 3) + 0.65y + 0.75z = 8.45 0.5z + 1.5 + 0.65y + 0.75z = 8.45 1.25z + 0.65y + 1.5 = 8.45 1.25z + 0.65y = 6.95

OpenStudy (aripotta):

now do that to the second equation

OpenStudy (anonymous):

ok thank you so much!!!!!

OpenStudy (aripotta):

we're not done yet though. that's just the first step

OpenStudy (aripotta):

do you know what to do from here?

OpenStudy (anonymous):

um actually no! sorry but i'm terrible at math!

OpenStudy (aripotta):

can you do the second equation?

OpenStudy (anonymous):

whats the answer for x?

OpenStudy (anonymous):

Is it 695

OpenStudy (aripotta):

three variable equations have a long process to solving :/ you have to substitute "z + 3" for x in the second equation (in my first post). once you do that, you have a two variable systems of equations and it's easier. but you have to do em in order

OpenStudy (anonymous):

so it would be(z+3) +y+z

OpenStudy (aripotta):

yea, z + 3 + y + z = 14, which is simplified to 2x + y = 11 so here's your new system of equations: 2x + y = 11 1.25z + 0.65y = 6.95

OpenStudy (anonymous):

ok now what

OpenStudy (aripotta):

solve for y in the first equation

OpenStudy (anonymous):

ok i'm lost! foreal.

OpenStudy (aripotta):

omg.

OpenStudy (anonymous):

i know i'm sorry

OpenStudy (aripotta):